# Institutionalizing Folk Theories of Algorithms: How Multi-Channel Networks (MCNs) Govern Algorithmic Labor in Chinese Live-Streaming Industry

### v1

@@ 3.2.1. Nine-month Ethnography in Beijing and Changsha
We conducted a nine-month ethnographic study (April–December 2023) within three MCNs (AB, QC, XH) based in Beijing and Changsha, actively participating in their live-streaming practices. These two cities were chosen because they represent distinct yet influential hubs in China’s live-streaming economy: Beijing, as the nation’s capital, is a center of digital entrepreneurship and policy discourse (Cunningham et al., 2019); Changsha, though a smaller city by comparison, has emerged as a leading cultural production hub that attracts a large number of rural migrant streamers due to its thriving live-streaming industry (Xiao et al., 2024a).
Over the course of our nine-month ethnographic fieldwork, we focused on how MCNs and live-streamers interact around platform algorithms. In the first three months, we took an open and exploratory approach to learn how streamers get started, what their daily routines look like, and how MCNs provide support or step in to guide their work. During this period, we also picked up basic streaming skills ourselves and built trust with both streamers and MCN staff. In the middle phase of our fieldwork, we started noticing how often both MCNs and streamers talked about “the algorithm.” The algorithm wasn’t just seen as background technology but became a key part of everyday decision-making. MCNs used it to set performance expectations, give advice, and explain success or failure. Streamers, in turn, developed their own informal beliefs about how the algorithm works, often based on MCN coaching. In the final four months, we shifted our attention fully to this process. We closely examined how MCNs interpret vague and unpredictable algorithmic signals and turn them into concrete rules and routines for streamers to follow. This shift led us to understand MCNs not just as business managers, but as algorithmic intermediaries, the organizations that translate platform uncertainty into structured labor expectations through shared stories, metrics, and coaching strategies.
We believe ethnography is particularly well suited for this research because it enables access to the tacit routines, emotional investments, and situated negotiations that structure algorithmic labor, the factors often overlooked in formal accounts or interviews. Long-term immersion allowed us to observe how such folk theories of algorithms are socially constructed and organizationally reinforced, not just individually imagined.
@@ 3.2.2. Interviews with Live-Streamers and MCNs Practitioners
Semi-structured interviews are widely recognized as a valuable complement to ethnographic fieldwork, offering opportunities to explore emergent themes in greater depth (Thach et al., 2023). Following our nine-month ethnography, we conducted semi-structured interviews in the end of our ethnography with 24 live-streamers and 13 MCN practitioners to further investigate how streamers perceive algorithmic control and how MCNs interpret, mediate, or reinforce these algorithmic dynamics.
@@
The interview protocol, developed based on prior research on live-streamers (Wohn and Freeman, 2020; Wang et al., 2019; Wu et al., 2023) and MCNs (Vonderau, 2016; Liang, 2022; Siciliano, 2023), then focused on several core areas: challenges associated with platform algorithms, streamers’ understanding and navigation of algorithmic systems, the forms of algorithmic guidance or control exerted by MCNs, and the infrastructural or interpretive support MCNs provide to facilitate compliance or success. For live-streamers, sample questions included: “How do you know when the algorithm is working in your favor or against you?”; “Has your MCN ever provided you with explanations or advice about the algorithm?”; and “What kind of support—technical, emotional, or interpretive—do you receive from your MCN in navigating platform challenges?” For MCN practitioners, we asked: “How do you stay informed about changes in the platform’s algorithmic rules?”; “Do you provide training or informal guidance to streamers on how to ‘play to the algorithm’?”; and “Would you describe your role for live-streamers regarding platform algorithms?”
@@
Each interview lasted approximately 60 to 90 minutes. Throughout, we encouraged participants to critically reflect on how platform algorithms shape their daily labor and how MCNs influence these experiences, not only as organizational actors but also as mediators of algorithmic knowledge and expectations.
@@ 3.3. Data Analysis
All interviews were conducted and recorded in Chinese to ensure participants could communicate comfortably and effectively. Three researchers subsequently transcribed and translated these recordings into English for analysis. We conducted a thematic analysis, combining both deductive and inductive coding methods. Deductive codes were informed by our theoretical framework and interview protocol, centering on constructs such as algorithmic governance, mediation, risk transfer, and infrastructural support. Inductive codes emerged through open reading of the transcripts and fieldnotes, capturing unanticipated practices, metaphors, tensions, and expressions from participants. We analyzed our ethnographic fieldnotes and interview transcripts together, rather than treating them as separate datasets. This integrated approach allowed us to triangulate across different forms of data, what participants said in interviews and what we observed in situ, yielding richer contextualization and interpretive depth  (Xiao et al., 2024b). Drawing from traditions in interpretive ethnography and qualitative HCI, we treated interviews not as isolated accounts but as situated reflections shaped by ongoing organizational and algorithmic conditions observed during fieldwork  (Seaver, 2017).
@@
We adopted a reflexive thematic analysis approach  (Kuo et al., 2023; Xiao et al., 2024b), which foregrounds the researchers’ interpretive labor and views knowledge production as iterative and situated. To begin, all three researchers independently coded a subset of transcripts and fieldnotes to generate initial codes. These were compared in regular team discussions, where we negotiated code definitions, added emergent categories, and refined our shared codebook. Rather than calculate inter-coder reliability, we emphasized interpretive alignment through dialogue, clarification, and documentation of divergent readings, following qualitative HCI best practices. After completing the open coding phase, we clustered related codes into higher-order conceptual categories. We used affinity diagramming (Haskins Lisle et al., 2020) for organizing codes, identifying cross-cutting themes, and mapping connections between concepts. Throughout the process, we wrote analytic memos to document interpretive decisions, surprising findings, and theoretical insights, maintaining a reflexive audit trail that anchored our evolving understanding.
3.4. Positionality Statements
Our research team brings diverse yet complementary perspectives shaped by years of engagement with the live-streaming ecosystem. Five researchers in our team have been involved in long-term research experience into platform labor and digital media economies. One team member has conducted extensive collaborative work with MCN companies, affording us privileged access to organizational practices and streamers’ networks while also requiring ongoing reflexivity about the power dynamics embedded in such relationships. Another two researchers are trained in critical political economy and labor studies, providing a structural and historically grounded lens to interpret data related to labor precarity. One member has conducted multiple ethnographic and interview-based studies on live-streaming cultures and creators’ self-presentation, contributing cultural and interactional sensitivity to our analysis. This range of backgrounds enriched our analytical process while also necessitating deliberate reflection on how our disciplinary commitments and field relationships shaped the way we framed questions, interpreted participants’ accounts, and conceptualized key themes.
4. Findings: Constructing Dual Folk Theories: How MCNs Make Opaque Algorithms Actionable for Streamers? (RQ1)
@@ 3.4. Positionality Statements
Our findings show that both MCNs and live-streamers operate under the same fundamental condition: algorithmic opacity. Live-streaming platforms rarely disclose how their recommendation systems function. As a result, both MCNs and streamers must navigate a system that is unpredictable, constantly shifting, and difficult to decipher. For streamers, this opacity often leads to frustration and powerlessness. They cannot understand why one stream gains traction while another goes unnoticed, or why their platform visibility drops without warning. In search of guidance, many turn to MCNs and hope these institutions have insight into how to improve their algorithmic exposure. However, MCNs themselves are not privileged insiders. Like the streamers they manage, they do not have access to the algorithm’s inner workings. They are also left guessing.
To make sense of this uncertainty, MCNs observe performance trends, compare outcomes across multiple streamers, and develop what we call folk algorithmic theories: informal, experience-based beliefs about how the algorithm behaves, what kinds of content it rewards, and how it might be influenced. These folk beliefs are speculative rather than technical. They are grounded in trial-and-error and pattern recognition, not in verified mechanisms. As many MCN interviewees in our study admitted, there is no guaranteed formula for algorithmic success, no matter how closely streamers follow the advice they are given.
In other words, MCNs are fully aware that their guidance cannot ensure that any individual streamer will succeed under the algorithm’s logic. As a result, MCNs maintain two distinct but interconnected versions of algorithmic theory. Internally, MCNs recognize that the algorithm is unstable, opaque, and ultimately uncontrollable. This understanding allows them to make business decisions cautiously, using flexible planning and risk management. Externally, however, they offer streamers simplified and more confident theories of algorithmic visibility. They portray the algorithm as fair, learnable, and responsive to effort. The message is clear: “If you keep trying, and follow our advice, you’ll succeed.” This theory of algorithms helps recruit and retain streamers, many of whom are willing to enter and stay in the industry, invest financially in equipment and services, and devote emotional energy in hopes of eventually gaining algorithmic visibility.
In the sections that follow, we show how MCNs carefully maintain these two theories—one inward-facing and probabilistic (subsection 4.1), the other outward-facing and prescriptive (subsection 4.2). Together, these two versions of algorithms theories allow MCNs to operate effectively in a deeply uncertain environment while preserving organizational authority and extracting economic value from precarious digital labor.
@@
4.1. Internal Folk Theories: Managing Uncertainty and Risk within MCNs
Despite their professional positioning and close partnership with platforms, MCNs do not possess privileged access to the internal workings of recommendation algorithms. For many MCN staff, the algorithm remains an unpredictable and opaque black box. “We don’t have backend access,” explained an operations manager (M3). “All our understanding of the algorithm is based on repeated trial and error (M10).” Faced with this uncertainty, MCNs develop their own internally shared folk theories within company staff to interpret platform behavior and improve streamers’ algorithmic visibility.
@@ 4.1. Internal Folk Theories: Managing Uncertainty and Risk within MCNs
These internal folk theories are expressed through heuristic rules and behavioral scripts, often referred to within the organizations as “formulas” or “track logics.” Typical examples include recommendations on streaming frequency (e.g., “at least three sessions per week”), minimum broadcast duration (e.g., “streams under 40 minutes won’t get recommended”) and emotional tone (e.g., “you must be excited in the first five minutes, or the system will judge it as low-interaction”). These behavioral prescriptions are formalized in onboarding materials, reiterated in weekly coaching sessions, and continuously updated via internal chat groups. Beyond behavioral scripts, many MCNs also develop aesthetic and infrastructural prescriptions that they believe enhance a streamer’s algorithmic visibility. For instance, staff often emphasize the algorithm’s supposed preference for high-definition visuals and stable audio quality, as well as the physical environment of streaming.
@@
However, while these theories offer a sense of control, most MCN staff readily admit their limitations. “Even if you follow all the best practices, sometimes the platform just doesn’t give you any algorithmic visibility,” M9 remarked. “And we don’t really know why.” This tension between the promise of algorithmic optimization and the reality of its unpredictability has prompted many MCNs to shift focus. Rather than attempting to fully decode the algorithm, MCNs increasingly prioritize monetization logic, developing what might be called revenue-first strategies. M9 explained, ”We’ve had many cases where someone did everything right—good lighting, good engagement, great voice—and still, no algorithmic visibility. There’s just no guarantee.”
@@
In response, MCNs adopt a statistical approach to managing uncertainty. Success is treated as a probability game: by recruiting large cohorts of new streamers and investing minimal resources in each, MCNs reduce the risk associated with any one individual’s failure. The goal is not to ensure the success of every recruit but to ensure that some will be algorithmically “picked.” As M9 elaborated: ”This process is highly random. We recruit newcomers who look attractive, provide them with some equipment, and ask them to start live-streaming. But we don’t know who will make it. What we do know is that if we have enough people, a certain percentage will succeed. Then we allocate resources to those who are ‘chosen’ by the algorithm.” This batch recruitment model allows MCNs to scale labor while minimizing upfront investment and maximizing return on attention. But the logic does not stop at internal risk distribution. Just as MCNs construct folk theories internally to manage uncertainty, they also construct simplified algorithmic theories externally—targeted at streamers.
@@
Unlike internal folk theories, which circulate among MCN staff as tentative, probabilistic tools for interpreting platform behavior, external folk theories are simplified and confident theories constructed for streamers. These externally projected theories serve two strategic purposes: first, they attract and retain labor by portraying algorithmic success as predictable and achievable, thereby encouraging newcomers to join and existing streamers to continue; second, they stimulate financial investment by persuading streamers to purchase services, equipment, or training packages under the belief that these will enhance their algorithmic performance, allowing MCNs to profit from these sales.
@@
While most algorithmic interpretations remain speculative for MCNs, certain practices appear to yield consistent, if not guaranteed, results. For example, some MCNs staffs (Such as M3, M4, M8, M9) report success in boosting algorithmic visibility by inputting large numbers of real-user accounts as followers for new streamers or by orchestrating coordinated entry of verified user accounts into live-stream rooms. These strategies are believed to trigger the algorithm’s visibility mechanisms, resulting in increased exposure and higher likelihood of recommendation. However, it is crucial to emphasize that even these seemingly “effective” tactics operate within a system of high variability and limited transparency. What works under one live-streaming platform version or for one streamer may not hold for another, reinforcing the broader uncertainty that underpins MCNs’ dual theorization of algorithms. Moreover, for most ordinary streamers, MCNs are often unwilling to invest significant resources into these tactics, given the high costs and uncertain returns. Thus, in this study, we do not seek to examine how MCNs develop sophisticated understandings of platform algorithms in order to manipulate them—indeed, in our fieldwork, MCNs rarely invested substantial effort in decoding or directly manipulating the algorithm, which they largely viewed as infeasible, unrewarding, and not worth the effort given the high uncertainty and low return. Rather, our focus lies in how MCNs construct and mobilize algorithmic folk theories—as organizational narratives—to manage algorithmic uncertainty, structure labor practices, and sustain control over streamers within the live-streaming ecosystem, a strategy that, in the context of our sampled Chinese MCNs, represents a more direct and prevalent path to profitability than algorithmic manipulation.
@@
Ultimately, what finally emerges is a dual system of epistemic practice about algorithmic folk theories: internally, MCNs recognize the algorithm as volatile and uncontrollable; externally, they reframe it as predictable and improvable. This asymmetry is not incidental—it is the foundation of the MCN business model. By selectively mobilizing folk theories in different contexts, MCNs manage precarity, sustain labor supply, and extend their authority within China’s live-streaming industry. Next, in subsection 4.2, we further analyze how these externally directed folk theories of algorithms are strategically constructed by MCNs for streamers.
4.2. External Folk Theories: Simplified Narrative to Manage and Govern Streamers

### v2

@@ 3.2.1. Nine-month Ethnography in Beijing and Changsha
We conducted a nine-month ethnographic study (April–December 2023) within three MCNs (A, B, C) based in Beijing and Changsha, actively participating in their live-streaming practices. These two cities were chosen because they represent distinct yet influential hubs in China’s live-streaming economy: Beijing, as the nation’s capital, is a center of digital entrepreneurship and policy discourse (Cunningham et al., 2019); Changsha, though a smaller city by comparison, has emerged as a leading cultural production hub that attracts a large number of rural migrant streamers due to its thriving live-streaming industry (Xiao et al., 2024a).
Over the course of our nine-month ethnographic fieldwork, we focused on how MCNs and live-streamers interact around platform algorithms. In the first three months, we took an open and exploratory approach to learn how streamers get started, what their daily routines look like, and how MCNs provide support or step in to guide their work. During this period, we also picked up basic streaming skills ourselves and built trust with both streamers and MCN staff. In the middle phase of our fieldwork, we started noticing how often both MCNs and streamers talked about “the algorithm.” The algorithm became a key part of everyday practice. MCNs used it to set performance expectations, give advice, and explain success or failure. Streamers, in turn, developed their own informal beliefs about how the algorithm works, often based on MCN coaching. In the final four months, we shifted our attention fully to this process. We closely examined how MCNs interpret vague and unpredictable algorithmic signals and turn them into concrete rules and routines for streamers to follow. This shift led us to understand MCNs both as business managers and as algorithmic intermediaries that translate platform uncertainty into structured labor expectations.
@@ 3.2.2. Interviews with Live-Streamers and MCNs Practitioners
Following our nine-month ethnography, we conducted semi-structured interviews in the end of our ethnography with 24 live-streamers and 13 MCN practitioners to further investigate how MCNs mediate algorithmic dynamics.
@@
The interview protocol, developed based on prior research on live-streamers (Wohn and Freeman, 2020; Wang et al., 2019; Wu et al., 2023; Chen et al., 2024) and MCNs (Vonderau, 2016; Liang, 2022; Siciliano, 2023), then focused on several core areas: challenges associated with platform algorithms, streamers’ understanding and navigation of algorithmic systems, the forms of algorithmic guidance or control exerted by MCNs, and the infrastructural or interpretive support MCNs provide to facilitate compliance or success. For live-streamers, sample questions included: “How do you know when the algorithm is working in your favor or against you?”; “Has your MCN ever provided you with explanations or advice about the algorithm?”; and “What kind of support—technical, emotional, or interpretive—do you receive from your MCN in navigating platform challenges?” For MCN practitioners, we asked: “How do you stay informed about changes in the platform’s algorithmic rules?”; “Do you provide training or informal guidance to streamers on how to ‘play to the algorithm’?”; and “Would you describe your role for live-streamers regarding platform algorithms?”
@@
Each interview lasted approximately 60 to 90 minutes. Throughout, we encouraged participants to critically reflect on how platform algorithms shape their everyday labor and how MCNs influence these experiences as organizational actors and mediators of algorithmic knowledge and expectations.
3.2.3. Interviews with Former Live-Streamers
To address potential survivor bias in studies of platform labor, after our ethnography and interviews with streamers and MCNs staffs, we also conducted semi-structured interviews with seven former live-streamers (F1–F7) who had previously been affiliated with MCNs and had exited the live-streaming industry (see Table 3). We realized accounts from active and relatively successful workers may overrepresent perspectives shaped by continued participation and organizational alignment, while obscuring experiences of disillusionment. Including former live-streamers enabled us to examine how MCN practices and algorithmic interpretations were reassessed after participants were no longer embedded in ongoing managerial relationships. Each interview lasted approximately 60 to 90 minutes.
Former live-streamers provided retrospective accounts that spanned both their period of active participation and their post-exit reflections. These interviews focused on how MCNs explained platform algorithms, how such explanations shaped streamers’ expectations and labor decisions, and how participants evaluated these narratives after leaving the industry. Sample questions included: “Looking back, how do you now interpret the advice and explanations provided by your MCN?”; and “What role did algorithm-related expectations and MCNs you met play in your decision to leave?”
3.2.4. Reflection of Study Design
Given the evaluative and hierarchical nature of MCN-managed live-streaming work, we remained attentive to how organizational conditions might shape participants’ accounts during ethnographic observation and interviews. We noticed that MCNs exercise varying degrees of control over streamers’ income, visibility, and access to future opportunities, which can influence how experiences are narrated in research settings. We also recognized that some MCN employees might be reluctant or hesitant to articulate critical reflections on organizational practices, particularly when such discussions could be perceived as professionally sensitive.
To ensure the reliability of our findings, we adopted several complementary methodological strategies. First, we emphasized confidentiality as well as the separation of research activities from organizational evaluation or performance assessment processes during interviews. We encouraged participants to begin with detailed descriptions of everyday work routines, which allowed accounts of algorithmic guidance and MCNs practices to emerge through situated examples.
Second, we conducted a relatively extensive set of interviews, which allowed recurring themes to emerge across accounts. Our interviewees included MCN staff in relatively senior positions like executive, who were less constrained by immediate power dependencies and more willing to speak candidly about organizational strategies. We also interviewed former streamers who no longer maintained contractual or financial ties with MCNs, reducing concerns about retaliation or reputational risk. These interviews provided perspectives that were less shaped by ongoing dependency relationships.
Third, we relied on systematic triangulation across multiple data sources. Interview data were examined in conjunction with long-term ethnographic observations, informal conversations, internal training materials, and accounts from MCN practitioners, streamers, and former streamers. The longitudinal nature of our ethnography also provided repeated opportunities to observe how algorithmic narratives were enacted in practice, reinforced through training and evaluation, and reproduced across everyday interactions.
@@ 3.3. Data Analysis
All interviews were conducted and recorded in Chinese to ensure participants could communicate comfortably and effectively. Three researchers subsequently transcribed and translated these recordings into English for analysis. We conducted a thematic analysis, combining both deductive and inductive coding methods. Deductive codes were informed by our theoretical framework and interview protocol, centering on constructs such as algorithmic governance, mediation, risk transfer, and infrastructural support. Inductive codes emerged through open reading of the transcripts and fieldnotes, capturing unanticipated practices, metaphors, tensions, and expressions from participants. We analyzed our ethnographic fieldnotes and interview transcripts together, rather than treating them as separate datasets. This integrated approach allowed us to triangulate across different forms of data, what participants said in interviews and what we observed in situ, yielding richer contextualization and interpretive depth  (Xiao et al., 2024b). Drawing from traditions in interpretive ethnography and qualitative HCI, we treated interviews as situated reflections shaped by ongoing organizational and algorithmic conditions observed during fieldwork  (Seaver, 2017).
@@
We adopted a reflexive thematic analysis (Braun and Clarke, 2019), following the traditional qualitative HCI approach  (Kuo et al., 2023; Xiao et al., 2024b, 2025), which focuses on the interpretive work of the researchers and views knowledge production as iterative and situated. To begin, all three researchers independently coded a subset of transcripts and fieldnotes to generate initial codes. These were compared in regular team discussions, where we negotiated code definitions, added emergent categories, and refined our shared codebook. Rather than calculate inter-coder reliability, we emphasized interpretive alignment through dialogue, clarification, and documentation of divergent readings, following qualitative HCI best practices. After completing the open coding phase, we clustered related codes into higher-order conceptual categories. We used affinity diagramming (Haskins Lisle et al., 2020) for organizing codes, identifying cross-cutting themes, and mapping connections between concepts.
3.4. Positionality Statement
Our research team brings situated perspectives shaped by years of engagement with China’s live-streaming ecosystem through research and industry experience. Our team members conducted extensive research on platform labor and digital economies, including several published paper on MCNs, live-streamers and platform policy. This background informed our sensitivity to everyday labor routines and the informal norms through which platform work is organized.
The research team’s engagement with MCN companies enabled access to organizational meetings, training sessions, and internal documents that would otherwise be difficult to observe. During the course of this engagement, one member of the research team contributed professional technical reports and analytical materials for multiple MCN companies, work that led to sustained collaboration and, at one point, an invitation to join an MCN as a formal employee. This experience provided firsthand exposure to internal processes, performance evaluation logics, and the organizational circulation of algorithm-related knowledge. Such role immersion are also characteristic of ethnography, in which researchers gain access and insight through sustained involvement in organizational practice (Dumont, 2023).
At the same time, such proximity required ongoing reflexive attention to power asymmetries and role boundaries, particularly when interpreting managerial narratives and performance-oriented accounts. Throughout the study, organizational perspectives were examined alongside ethnographic observations and interviews with active and former streamers. This analytic triangulation helped ensure that institutional viewpoints were situated within a broader field of perspectives and did not dominate the interpretation of algorithmic practices and labor relations. We engaged in regular reflexive discussions to examine how prior field relationships and assumptions about platform work shaped our research. These discussions informed the refinement of interview protocols and the articulation of key analytical concepts.
4. Findings: MCNs’ Discursive Construction of Algorithms (RQ1)
@@
Live-streaming platforms rarely disclose how their recommendation systems function. As a result, live-streamers must navigate a system that is unpredictable, constantly shifting, and difficult to decipher. For streamers, this opacity often leads to frustration and powerlessness. They cannot understand why one stream gains traction while another goes unnoticed, or why their platform visibility drops without warning. This uncertainty is particularly acute for streamers with limited formal education or prior work experience, a common profile within the live-streaming workforce. Lacking alternative sources of technical or professional guidance, these streamers often rely heavily on MCNs’ explanations of how algorithms work and how exposure can be improved. MCNs thus become a primary interpretive authority through which algorithmic outcomes are understood.
At the same time, during our ethnography and interviews, we found that MCNs themselves do not occupy a privileged insider position. Like the streamers they manage, MCNs also lack access to the internal mechanism of platform algorithms and must operate through inference, observation, and accumulated experience. To make sense of this algorithmic uncertainty, MCNs observe performance trends, compare outcomes across multiple streamers, and develop related algorithmic “knowledge”: informal, experience-based beliefs about how the algorithm behaves, what kinds of content it rewards, and how it might be influenced. They are grounded in trial-and-error and pattern recognition, not in verified mechanisms. Guidance derived from these beliefs is widely understood within MCNs as having the potential to improve algorithmic visibility at an aggregate level. MCN practitioners often described such advice as “generally helpful” (M6) or “likely to work for most streamers” (M7), while acknowledging its limits in practice. They could not determine whether a given recommendation would be effective for any specific streamer, nor could they specify the magnitude or durability of its effects. As many MCN interviewees in our study admitted, there is no formula that ensures algorithmic success, even when streamers closely follow the advice they are given. Despite this internal shared uncertainty, MCNs often present their algorithmic explanations with confidence and clarity when communicating with streamers and instructing streamers how to deal with algorithms. This asymmetry between internal uncertainty and external certainty forms the basis for how MCNs construct and disseminate algorithmic interpretations within live-streaming industry. We thus focus on how MCNs intend to construct their algorithmic interpretations in this paper.
Here, we decide to use the term discursive construction of algorithms to characterize MCNs’ algorithm-related explanations because these accounts function as intentional and strategic interventions to streamers’ behavior. Unlike folk theories, which typically emerge as users’ folk attempts to understand opaque systems, MCNs’ algorithm explanations are developed and deployed at organizational level. They are crafted to coordinate behavior across multiple streamers, standardize content production practices, and justify managerial directives. Algorithmic explanations become tools for governing labor rather than personal interpretations held by individual users. Similarly, while the concept of algorithmic imaginaries emphasizes how users envision algorithms and their social implications, MCNs’ practices extend beyond imagination. Their discursive constructions selectively amplify certain algorithmic attributes, such as rewardability, while downplaying uncertainty. This selective emphasis serves concrete organizational goals, particularly the management of streamer labor and the pursuit of increasing profitability.
The discursive construction of algorithms by MCNs includes internal and external interpretations to manage labor and increase profits. In the sections that follow, we show how MCNs construct these two interpretations—one inward-facing and probabilistic (subsection 4.1), the other outward-facing and prescriptive (subsection 4.2). This section 4 here focuses on the substance of these interpretations, that is, what MCN staff and streamers come to believe about how the algorithm works within this explanatory framework. In following section 5, we extend this analysis by examining how these algorithmic interpretations are mobilized in labor management practices.
@@
4.1. Internal Algorithmic Interpretation: Managing Uncertainty and Risk within MCNs
As mentioned by M1, the executive of a famous MCN, MCN’s algorithmic understandings are initially hoped to be oriented toward increasing the algorithmic visibility of affiliated streamers, thereby generating higher traffic, longer viewing times, and ultimately greater live-streaming revenue. Improved streamer performance translates into higher commission income, stronger bargaining power with platforms, and greater profitability for MCNs themselves. To achieve this goal, MCNs accumulated operational experience and the circulation of informal industry knowledge. Staff continuously test different content formats, streaming schedules, and interaction styles across multiple streamers, while also exchanging observations through personal networks, platform rumors, and peer organizations. These diverse sources are synthesized into actionable understandings of how algorithmic visibility might be advanced under conditions of opacity.
@@ 4.1. Internal Algorithmic Interpretation: Managing Uncertainty and Risk within MCNs
These internal understandings of algorithms are expressed through heuristic rules and behavioral scripts, often referred to within organizations as “formulas” or “track logics.” Typical examples include recommendations on streaming frequency (e.g., “at least three sessions per week”), minimum broadcast duration (e.g., “streams under 40 minutes won’t get recommended by algorithms”), and emotional tone (e.g., “you must be excited in the first five minutes, or the system will judge it as low-interaction”). Beyond behavioral scripts, many MCNs also develop aesthetic and infrastructural prescriptions that they believe enhance algorithmic visibility. Staff frequently emphasize the importance of high-definition visuals, stable audio quality, lighting conditions, and even the physical arrangement of the streaming environment. These elements are framed as signals that the algorithm is assumed to recognize and reward.
@@
However, in practice, while these algorithmic interpretations offer a sense of control, most MCN staff readily admit their limitations, no matter how much efforts MCNs made to understand algorithms. “Even if you follow all the best practices, sometimes the platform just doesn’t give you any algorithmic visibility,” M9 remarked. This tension between the promise of algorithmic optimization and the reality of its unpredictability has prompted many MCNs to shift focus. Rather than attempting to fully decode the algorithm, MCNs increasingly prioritize monetization logic.
@@
MCNs continue to hope to increase the visibility of each affiliated streamer and to provide support toward that goal. At the same time, they are fully aware of the algorithm’s fundamental unpredictability. As a result, MCNs adopt a statistical approach to profitability that treats uncertainty as something to be managed rather than eliminated. By recruiting large cohorts of new streamers and investing minimal resources in each, MCNs reduce the risk associated with any one individual’s failure. The goal is not to ensure the success of every recruit but to ensure that some will be algorithmically “picked.” As M9 elaborated: “This process is highly random. We recruit newcomers who look attractive, provide them with some equipment, and ask them to start live-streaming. But we don’t know who will make it. What we do know is that if we have enough people, a certain percentage will succeed. Then we allocate resources to those who are ‘chosen’ by the algorithm.” This batch recruitment model allows MCNs to scale labor while minimizing upfront investment and maximizing return on attention.
@@
But the logic does not stop at internal risk distribution. Just as MCNs construct algorithmic interpretation internally to manage uncertainty, they also construct simplified algorithmic interpretation externally—targeted at streamers. Unlike internal interpretation, which circulate among MCN staff as tentative, probabilistic tools for interpreting platform behavior, external interpretations are simplified and confident theories constructed for streamers. These externally projected interpretations serve two strategic purposes: first, they attract and retain labor by portraying algorithmic success as predictable and achievable, thereby encouraging newcomers to join and existing streamers to continue; second, they stimulate financial investment by persuading streamers to purchase services, equipment, or training packages under the belief that these will enhance their algorithmic performance, allowing MCNs to profit from these sales.
@@
In this study, we do not seek to fully examine how MCNs develop sophisticated understandings of platform algorithms in order to manipulate them—indeed. In our fieldwork, MCNs rarely invested substantial effort in decoding or directly manipulating the algorithm, which they largely viewed as infeasible, unrewarding, and not worth the effort given the high uncertainty and low return. Rather, our focus lies in how MCNs construct and mobilize discursive construction of algorithms—as organizational narratives—to manage algorithmic uncertainty, structure labor practices, and sustain control over streamers within the live-streaming ecosystem. This is a strategy that, in the context of our sampled Chinese MCNs, represents a more direct and prevalent path to profitability than algorithmic manipulation.
@@
What finally emerges is a dual system of epistemic practice about algorithmic interpretations: internally, MCNs recognize the algorithm as volatile and uncontrollable; externally, they reframe it as predictable and improvable. This asymmetry is not incidental—it is the foundation of the MCN business model. By selectively mobilizing algorithmic interpretations in different contexts, MCNs manage precarity, sustain labor supply, and extend their authority within China’s live-streaming industry. Next, in subsection 4.2, we further analyze how these externally directed algorithmic interpretations are strategically constructed by MCNs for streamers.
4.2. External Interpretation: Simplified Narrative to Manage and Govern Streamers